### v1

@@ Abstract
Neural operators learn mappings between function spaces, which is practical for learning solution operators of PDEs and other scientific modeling applications. Among them, the Fourier neural operator (FNO) is a popular architecture that performs global convolutions in the Fourier space. However, such global operations are often prone to over-smoothing and may fail to capture local details. In contrast, convolutional neural networks (CNN) can capture local features but are limited to training and inference at a single resolution. In this work, we present a principled approach to operator learning that can capture local features under two frameworks by learning differential operators and integral operators with locally supported kernels. Specifically, inspired by stencil methods, we prove that we obtain differential operators under an appropriate scaling of the kernel values of CNNs. To obtain local integral operators, we utilize suitable basis representations for the kernels based on discrete-continuous convolutions. Both these approaches preserve the properties of operator learning and, hence, the ability to predict at any resolution. Adding our layers to FNOs significantly improves their performance, reducing the relative $L^{2}$-error by $34$-$72$% in our experiments on turbulent 2D Navier-Stokes fluid flow and the spherical shallow water equations.
@@ 1 Introduction
Deep learning holds the promise to greatly accelerate advances in computational science and engineering, which often require numerical solutions of partial differential equations (PDEs) Azzizadenesheli et al. (2023); Zhang et al. (2023); Cuomo et al. (2022). Recent advances in deep learning have enabled applications such as weather forecasting (Pathak et al., 2022; Bonev et al., 2023; Lam et al., 2023), seismology (Sun et al., 2023; Shi et al., 2023), reservoir engineering for carbon capture (Wen et al., 2022; 2023), and many other applications with orders of magnitude speedup over traditional methods.
@@
Many of the above results are achieved by neural operators, which learn mappings between function spaces, enabling operator learning for function-valued data (Li et al., 2021; Azzizadenesheli et al., 2023; Raonić et al., 2023). In particular, they are agnostic to the discretization of the input and output functions—a vital property in the context of PDEs where data is often provided at varying resolutions and high-resolution data is costly to generate Kovachki et al. (2021). In contrast, standard neural networks such as convolutional neural networks (CNN) (Ronneberger et al., 2015; Gupta & Brandstetter, 2022) require the functions to be discretized at a fixed resolution on a regular grid, which is limiting.
@@
In the past few years, various architectures of neural operators have been developed. Among them, the Fourier neural operator (FNO) Li et al. (2020a) has gained popularity and shown good performance in a number of applications. FNO performs global convolutions in the Fourier space. However, such global operations are often prone to over-smoothing and may fail to capture local details.
@@
There are many applications that require a local neural operator. For instance, solution operators of several relevant PDEs are of local nature. Examples include hyperbolic PDEs, which have real-valued characteristic curves LeVeque (1992). As a result, a solution at a given point will only depend on the initial condition within a neighborhood of that point. As such, their solution operators only have a local receptive field and can, therefore, be efficiently learned by locally supported kernels.
@@
Further examples of local operators are differential operators, which can be expressed in terms of pointwise multiplication with the frequency in the spectral domain. Consequently, they introduce large errors when approximated by a finite number of parameters in Fourier space. In this context, we note that the emulation of classical numerical methods to solve partial differential equations, such as finite-difference methods, relies on the usage of local stencils for differentiation Thomas (2013). This calls for the presence of local and differential operators in neural operator architectures.
@@
Table 1: Comparison of different architectures for the solution of PDEs. The top half enumerates architectures for planar domains and the bottom half for spherical domains. Our proposed architectures are highlighted in bold letters, and the differential and integral kernels are detailed in Section 3.
@@
Naturally, every local operation can also be represented by a global operation. However, this is typically vastly parameter-inefficient and does not provide a good inductive bias for learning local operations. In the context of neural operators, spectral variants of neural operators, such as the FNO (Li et al., 2020a) and Spherical FNO (SFNO) (Bonev et al., 2023), are theoretically able to approximate local convolutions. However, representing local kernels requires the approximation of a global signal in the spectral domain, in turn demanding a large number of parameters (due to the uncertainty principle Cohen-Tannoudji et al. (1977)).
A few prior works have explored the usage of local operations in the context of neural operators. Li et al. (2020b) introduce graph neural operators (GNOs), which parameterize local integral kernels with a neural network. However, evaluating this network on all combinations of points to compute the integrals can make GNOs computationally expensive and slower than hardware-optimized convolutional kernels. Alternatively, Ye et al. (2022; 2023) propose local neural operators by combining FNOs with convolutional layers and, similarly, Wen et al. (2022) integrate U-Nets and FNOs. Moreover, Raonić et al. (2023) propose convolutional neural operators (CNOs) by leveraging U-Nets that (approximately) respect the bandlimits.
However, since all of the above approaches rely on standard convolutional layers on equidistant grids, they have the following shortcomings. First, such approaches do not allow for a natural extension to unstructured grids or other geometries, which are ubiquitous in PDE problems Li et al. (2023). Moreover, they can only be applied to higher resolutions by downsampling of the (intermediate) inputs to the training resolution. Important high-frequency content can be lost through downsampling, which is particularly problematic for multi-scale data in the context of PDEs. In contrast, we develop convolutional layers that can be applied at any resolution without downsampling. We achieve this by appropriately scaling the receptive field or the values of the kernel, see Figure 1.
To summarize, current neural operator architectures suffer from at least one of the following limitations (see Table 1): (1) they cannot succinctly represent operations with a local receptive field, e.g. FNO, or (2) they cannot be applied to different resolutions without relying on explicit up-/downsampling which may degrade performance, e.g., CNO, or (3) they cannot be scaled to obtain sufficient expressivity, since they incur prohibitively high computationally costs, such as, for instance, GNOs.
Regular kernel
Differential kernel
Local integral kernel
[…]
Figure 1: Visualization of different limits of a convolution with a discretized function $v$ as the grid width $h$ is refined, i.e., $h\to 0$. (Top) A regular convolution is collapsing to a pointwise linear operator. (Middle) When we rescale and constrain the kernel appropriately, we converge to a differential operator, see Section 3.2. (Bottom) Alternatively, we can use a kernel that can be evaluated at arbitrary resolutions and keep the receptive field unchanged to converge to a local integral operator, see Section 3.4.
@@
For the first, we draw inspiration from stencils of finite-difference methods. We derive conditions to modify convolutional layers such that they converge to a unique differential operator when the discretization is refined. For the second case of local integral operators, we adapt discrete-continuous (DISCO) convolutions Ocampo et al. (2022) to provide an efficient, discretization-agnostic framework that can be applied to general meshes on both planar and spherical geometries.
Finally, we devise efficient implementations of both layers and show that the inductive bias of local operations can significantly improve the performance of FNOs and SFNOs on three different benchmarks. In particular, we learn a differential operator in a Darcy flow setting for motivating the need for neural operators with local inductive biases, and we improve over FNO by $87\%$. Additionally, we improve over the baseline (S)FNO by $34\%$ on turbulent 2D Navier-Stokes equations and by $72\%$ on the shallow water equations on the sphere.
@@ 2 Related work and connections to other frameworks
We strive to formulate local operations that are consistent with the neural operator paradigm, which stipulates that operations within the model are independent of the discretization of input functions Kovachki et al. (2021). In principle, this allows the evaluation of the neural operator on arbitrary meshes [^1] As long as they are suitable for the evaluation of the numerical operations of the neural operator, such as the Discrete Fourier Transform used in the Fourier neural operator, see Section 3.. Several discretization-independent local operations have been introduced in the deep learning literature. We outline the connections between these works in the following.
@@
In general, local operations can be provided by graph neural operators (GNOs) Li et al. (2020b). However, in their general form, GNOs are typically slow, hard-to-train, and not equivariant w.r.t. the symmetry group of the underlying domain Li et al. (2020c). Taking a convolutional kernel, GNOs subsume a series of works on hypernetwork approaches for convolutional layers Wang et al. (2018); Shocher et al. (2020).
@@
We show that DISCO convolutions Ocampo et al. (2022) represent special cases of GNOs with a convolutional kernel, which enables generalization to different geometries, and an efficient implementation, since the kernel can be pre-computed (see Section 3). In particular, DISCO convolutions are typically implemented as learnable linear combinations of fixed basis functions instead of neural networks. Moreover, they satisfy the equivariance in the limit of exact integration for special function classes.
@@
In the area of computer vision, there has been a series of adaptations of CNNs to be (locally) scale-equivariant by using filter dilation, filter rescalings in the discrete domain Rahman & Yeh (2023); Sosnovik et al. (2021); Worrall & Welling (2019) or the continuous domain Xu et al. (2014); Sosnovik et al. (2019); Ghosh & Gupta (2019), or input rescalings Marcos et al. (2018). Moreover, filter rescaling has also been explored for more general group-convolutions Bekkers (2019).
@@
In combination with FNOs, interpolation of the input and output functions has been explored by Wen et al. (2022). Also Ye et al. (2023; 2022) emphasized the need for local convolutions in FNOs, however, they do not discuss the application to different discretizations. Combined with correct treatment of the bandlimit of the functions (based on Karras et al. (2021)), the convolutional neural operator Raonić et al. (2023) also uses up- and downsampling of the input and output functions to apply a U-Net architecture.
@@ 3.1 Motivation: Convolutional Layer
$z_{i}=h\left(i-1-\frac{S-1}{2}\right),$ (2)
@@ 3.2 Differential Layer
In this section, we construct a layer that converges to a differential operator when the width $h$ of the discretization $D^{h}$ tends to zero. To prevent the operator from collapsing to a pointwise operation, we constrain and rescale the values of the kernel (according to the discretization width). Note that without rescaling, we would again recover a pointwise operator; however, due to the rescaling by a factor of $\frac{1}{h}$, the limit might not exist without constraining the values of the kernel. The next proposition shows that it is sufficient to subtract the average kernel value.
@@
Proof.
See Appendix A for a more general statement and a corresponding proof. ∎
Proposition 3.1 shows that we can learn different directional derivatives using an appropriate adaptation of standard convolutional layers as in Section 3.1. Specifically, we center the kernel $K$ by subtracting its mean $\bar{K}$ and scale the result by the reciprocal resolution $\frac{1}{h}$.
@@ 3.3 Integral Kernel Layers
Instead of rescaling the kernel, we can also adapt the size of the kernel such that the receptive field stays the same, i.e., is independent of the resolution $h$. We will explain in the following how we can adapt the size of the kernel with a fixed number of parameters.
Graph neural operator (GNO).
@@
Fourier Neural Operator (FNO).
@@
The remaining question is centered around the parametrization of the kernel. One could draw inspiration from works on hypernetworks in computer vision and parameterize it using a neural network Shocher et al. (2020). However, this is significantly more costly than a standard convolution. Another idea is to interpolate a fixed-sized kernel, e.g., using sinc or bi-linear interpolation. A more general version of the latter, based on a learnable linear combination of hat functions, is also known as discrete-continuous (DISCO) convolution and has shown to be effective in different applications Ocampo et al. (2022). We note that this formulation is closest to the original convolutional layer typically used in computer vision. However, in the next section, we see that it also allows us to use unstructured meshes and formulate the operation on more general domains $D\subset\mathbb{R}^{d}$.
@@ 3.4 General discrete-continuous convolutions
where $g,x\in G$ and $\mathrm{d}\mu(x)$ is the invariant Haar measure on $G$. This formulation presents us with the challenge of being non-trivial to discretize. The framework of DISCO convolutions Ocampo et al. (2022) achieves this by approximating the integral with a quadrature rule while evaluating the group action continuously:
@@ 3.5 Local neural operator architecture
To design our neural operator, we want to combine pointwise, local, and global operations. To this end, we take an FNO (or SFNO for spherical problems) as a starting point, which already features global operations in Fourier space and pointwise operations using its residual connections. Then, we augment the operators by adding our proposed local convolutions from the previous section as additional branches in the respective layers (see Figure 2).
@@ 4 Experiments
To validate the effectiveness of local operators, we evaluate the architecture on three PDE problems. Figure 3 shows samples on two of these PDEs. In all three cases, we incorporate our proposed local operators into existing FNO and SFNO architectures to demonstrate that a significant improvement in performance can be achieved by introducing the inductive bias of local convolutions. This section describes the experimental setting for each problem. Specific implementation details [^5] Code will be made available in the neuraloperator library at https://github.com/neuraloperator/neuraloperator. and hyperparameters can be found in Appendix C.
4.1 Darcy flow
$-\nabla\cdot(a\nabla u)=f,\quad u|_{\partial D}=0,$ (10)
$u\mapsto-\nabla\cdot(a\nabla u).$ (11)
Following the setup of Hasani & Ward (2024), we take
$a(x)=\begin{bmatrix}x_{1}^{2}&\sin(x_{1}x_{2})\\ x_{1}+x_{2}&x_{2}\end{bmatrix}$
and
$u(x)=\sum_{i,j=1}^{20}\frac{c_{ij}}{\sqrt{\lambda_{ij}}}\sin(i\pi x_{1})\sin(j\pi x_{2}),$
for $x\in D$, where,
$\lambda_{ij}$ $=(i\pi)^{2}+(j\pi)^{2}$
$c_{ij}$ $\sim\mathcal{N}(0,1/(i+j)).$
We then compare the performance of our proposed models: FNO in parallel with differential kernels, FNO with integral kernels, and FNO with both kernels. We also compare with a baseline FNO and the U-Net architecture from Gupta & Brandstetter (2022). Moreover, we perform zero-shot super-resolution on this problem; these results can be found in Section C.4.
4.2 Navier-Stokes equations
$\nabla\cdot u$ $=0$
on the spatial domain $D=[0,2\pi]^{2}$ equipped with periodic boundary conditions. In the above, $u$ and $p$ denote the velocity and pressure, and Re denotes the Reynolds number of the flow. We want to learn the solution operator mapping initial conditions $u(\cdot,0)=u_{0}$ to the time-evolved solution in vorticity form $w(\cdot,\tau)$ at time $\tau\in(0,\infty)$, where the vorticity is given by
$w=(\nabla\times u)\hat{z}$ (13)
with $\hat{z}$ being a unit vector normal to the plane.
We use the dataset from Li et al. (2022), which sets $m=4$ and $\textrm{Re}=5000$ in (12) and uses a temporal discretization of $\tau=1$ on a $128\times 128$ regular grid. The initial conditions are sampled from a Gaussian measure as described in Li et al. (2022), and the equation is solved with a pseudo-spectral solver. We emphasize that learning the solution operator for such a high Reynolds number is very challenging due to the turbulent nature and small-scale features of the flow. We conjecture that the baseline FNO is prone to over-smoothing over the finer scales, consequently leading to a degradation of performance.
4.3 Shallow water equations
Finally, we test our proposed architecture on the shallow water equations on the rotating sphere. They represent a system of hyperbolic partial differential equations used to model a variety of geophysical flow phenomena, such as atmospheric flows, tsunamis, and storm surges. They can be formulated in terms of the evolution of two state variables $\varphi$ and $u$ (geopotential height and the tangential velocity of the fluid column), governed by the equations
[…]
$F_{ij}=\varphi u_{i}u_{j}+\frac{1}{2}\varphi^{2},$ (15)
and a source term
$f(x)=-2\Omega x\times(\varphi u),\quad x\in\mathbb{S}^{2},$ (16)
which models the Coriolis force due to the rotation of the sphere with angular velocity $\Omega$. We use the dataset from Bonev et al. (2023), which uses a Gaussian random field to generate initial conditions on the sphere at a resolution of $256\times 512$ and solves for $\varphi$, $u$ at a lead time of one hour. The target solution is computed using a spectral solver, which takes 24 Euler steps [^6] Dataset and solver are taken from the torch-harmonics package at https://github.com/NVIDIA/torch-harmonics..
As baselines, we use a planar U-Net architecture, a spherical U-Net, where convolutions are performed using local integral kernels on the sphere, and an SFNO architecture. On the sphere, we only consider the impact of the integral kernel on the SFNO architecture, as it allows for a natural extension to the sphere, see Section 3.4.

### v2

@@ Abstract
Neural operators learn mappings between function spaces, which is practical for learning solution operators of PDEs and other scientific modeling applications. Among them, the Fourier neural operator (FNO) is a popular architecture that performs global convolutions in the Fourier space. However, such global operations are often prone to over-smoothing and may fail to capture local details. In contrast, convolutional neural networks (CNN) can capture local features but are limited to training and inference at a single resolution. In this work, we present a principled approach to operator learning that can capture local features under two frameworks by learning differential operators and integral operators with locally supported kernels. Specifically, inspired by stencil methods, we prove that we obtain differential operators under an appropriate scaling of the kernel values of CNNs. To obtain local integral operators, we utilize suitable basis representations for the kernels based on discrete-continuous convolutions. Both these approaches preserve the properties of operator learning and, hence, the ability to predict at any resolution. Adding our layers to FNOs significantly improves their performance, reducing the relative $L^{2}$-error by $34$-$72$% in our experiments, which include a turbulent 2D Navier-Stokes and the spherical shallow water equations.
Neural Operators, Scientific Machine Learning, Convolutions
@@ 1 Introduction
Regular kernel
Local integral kernel
Differential kernel
[…]
Figure 1: Visualization of different limits of a convolution with a discretized function $v$ as the grid width $h$ is refined, i.e., $h\to 0$. (Top) A regular convolution is collapsing to a pointwise linear operator. (Middle) Instead, we can use a kernel that can be evaluated at arbitrary resolutions and keep the receptive field unchanged, to converge to a local integral operator, see Section 3.4. (Bottom) Alternatively, we can let it collapse while constraining the kernel appropriately, converging to a differential operator, see Section 3.2.
@@
Deep learning holds the promise to greatly accelerate advances in computational science and engineering, which often require numerical solutions of partial differential equations (PDEs) (Azizzadenesheli et al., 2024; Zhang et al., 2023; Cuomo et al., 2022). Recent advances in deep learning have enabled applications such as weather forecasting (Pathak et al., 2022; Bonev et al., 2023; Lam et al., 2023), seismology (Sun et al., 2023; Shi et al., 2023), reservoir engineering for carbon capture (Wen et al., 2022, 2023), and many other applications with orders of magnitude speedup over traditional methods.
@@
Many of the above results are achieved by neural operators, which learn mappings between function spaces, enabling operator learning for function-valued data (Li et al., 2021; Azizzadenesheli et al., 2024; Raonić et al., 2023). In particular, they are agnostic to the discretization of the input and output functions—a vital property in the context of PDEs where data is often provided at varying resolutions and high-resolution data is costly to generate (Kovachki et al., 2021). In contrast, standard neural networks such as convolutional neural networks (CNN) (Ronneberger et al., 2015; Gupta & Brandstetter, 2022) require the functions to be discretized at a fixed resolution on a regular grid, which is limiting.
@@
In the past few years, various architectures of neural operators have been developed. Among them, the Fourier neural operator (FNO) (Li et al., 2020a), which performs global convolutions in Fourier space, has gained popularity and shown good performance in a number of applications. However, such global operations are often prone to over-smoothing and may fail to capture local details. Several other architectures instead learn global operators in physical space (e.g., (Li et al., 2022b)), but they likewise lack the inductive bias of local operations.
@@
There are many applications that require a local neural operator. For instance, solution operators of several relevant PDEs are of local nature. Examples include hyperbolic PDEs, which have real-valued characteristic curves (LeVeque, 1992). As a result, a solution at a given point will only depend on the initial condition within a neighborhood of that point. As such, their solution operators only have a local receptive field and can, therefore, be efficiently learned by locally supported kernels.
@@
Further examples of local operators are differential operators, which can be expressed in terms of pointwise multiplication with the frequency in the spectral domain. Consequently, they introduce large errors when approximated by a finite number of parameters in Fourier space. In this context, we note that the emulation of classical numerical methods to solve partial differential equations, such as finite-difference methods, relies on the usage of local stencils for differentiation (Thomas, 2013). This calls for the presence of local and differential operators in neural operator architectures.
@@
Table 1: Comparison of different architectures for the solution of PDEs. The top half enumerates architectures for planar domains and the bottom half for spherical domains. Our proposed architectures are highlighted in bold and the differential and integral kernels are detailed in Section 3.
@@
Naturally, every local operation can also be represented by a global operation. However, this is typically vastly parameter-inefficient and does not provide a good inductive bias for learning local operations. In the context of neural operators, spectral variants of neural operators, such as the FNO (Li et al., 2020a) and Spherical FNO (SFNO) (Bonev et al., 2023), are theoretically able to approximate local convolutions. However, representing local kernels requires the approximation of a global signal in the spectral domain, in turn demanding a large number of parameters (due to the uncertainty principle (Cohen-Tannoudji et al., 1977)).
A few prior works have explored the usage of local operations in the context of neural operators. Li et al. (2020b) introduce graph neural operators (GNOs), which parameterize local integral kernels with a neural network. However, evaluating this network on all combinations of points to compute the integrals can make GNOs computationally expensive and slower than hardware-optimized convolutional kernels. Alternatively, Ye et al. (2022, 2023) propose local neural operators by combining FNOs with convolutional layers and, similarly, Wen et al. (2022) integrate U-Nets and FNOs. Moreover, Raonić et al. (2023) propose convolutional neural operators (CNOs) by leveraging U-Nets that (approximately) respect the bandlimits.
However, since all of the above approaches rely on standard convolutional layers on equidistant grids, they have the following shortcomings. First, such approaches do not allow for a natural extension to unstructured grids or other geometries, which are ubiquitous in PDE problems (Li et al., 2023). Moreover, they can only be applied to higher resolutions by downsampling of the (intermediate) inputs to the training resolution. Important high-frequency content can be lost through downsampling, which is particularly problematic for multi-scale data in the context of PDEs. In contrast, we develop convolutional layers that can be applied at any resolution without downsampling. We achieve this by appropriately scaling the receptive field or the values of the kernel (see Figure 1).
To summarize, current neural operator architectures suffer from at least one of the following limitations (see Table 1): (1) they cannot succinctly represent operations with a local receptive field, e.g., FNO, or (2) they cannot be applied to different resolutions without relying on explicit up-/downsampling which may degrade performance, e.g., CNO, or (3) they cannot be scaled to obtain sufficient expressivity, since they incur prohibitively high computationally costs, such as, for instance, GNOs.
@@
For the first, we draw inspiration from stencils of finite-difference methods. We derive conditions to modify convolutional layers such that they converge to a unique differential operator when the discretization is refined. For the second case of local integral operators, we adapt discrete-continuous (DISCO) convolutions (Ocampo et al., 2022) to provide an efficient, discretization-agnostic framework that can be applied to general meshes on both planar and spherical geometries. To our knowledge, we are the first to connect the DISCO framework to operator learning.
Finally, we devise efficient implementations of both layers and show that the inductive bias of local operations (see Figure 2) can significantly improve the performance of FNOs and SFNOs on three different benchmarks. In particular, we learn a differential operator in a Darcy flow setting for motivating the need for neural operators with local inductive biases, and we improve over FNO by $87\%$. Additionally, we improve over the baseline (S)FNO by $34\%$ on turbulent 2D Navier-Stokes equations, by $72\%$ on the shallow water equations on the sphere, and by $63\%$ on the 2D diffusion-reaction equation. We also apply our method on modeling a flow past a cylinder, discretized on an irregular grid. We improve upon the best baseline by $42\%$.
@@ 2 Related work and connections to other frameworks
We strive to formulate local operations that are consistent with the neural operator paradigm, which stipulates that operations within the model are independent of the discretization of input functions (Kovachki et al., 2021). In principle, this allows the evaluation of the neural operator on arbitrary meshes [^1] As long as they are suitable for the evaluation of the numerical operations of the neural operator, such as the Discrete Fourier Transform used in the Fourier neural operator (see Section 3).. Several discretization-independent local operations have been introduced in the deep learning literature. We outline the connections between these works in the following.
@@
In general, local operations can be provided by graph neural operators (GNOs) (Li et al., 2020b). However, in their general form, GNOs are typically slow, hard-to-train, and not equivariant w.r.t. the symmetry group of the underlying domain (Li et al., 2020c). Taking a convolutional kernel, GNOs subsume a series of works on hypernetwork approaches for convolutional layers (Wang et al., 2018; Shocher et al., 2020).
@@
We show that DISCO convolutions (Ocampo et al., 2022) represent special cases of GNOs with a convolutional kernel, which enables generalization to different geometries and an efficient implementation, since the kernel can be pre-computed (see Section 3). In particular, DISCO convolutions are typically implemented as learnable linear combinations of fixed basis functions instead of neural networks. Also, they define maps that are equivariant up to the error resulting from the quadrature rule.
@@
In the area of computer vision, there has been a series of adaptations of CNNs to be (locally) scale-equivariant by using filter dilation, filter rescalings in the discrete domain (Rahman & Yeh, 2023; Sosnovik et al., 2021; Worrall & Welling, 2019) or the continuous domain (Xu et al., 2014; Sosnovik et al., 2019; Ghosh & Gupta, 2019), or input rescalings (Marcos et al., 2018). Moreover, filter rescaling has also been explored for more general group-convolutions (Bekkers, 2019).
@@
In combination with FNOs, interpolation of the input and output functions has been explored by Wen et al. (2022). Also, Ye et al. (2023, 2022) emphasized the need for local convolutions in FNOs, however, they do not discuss the application to different discretizations. Combined with correct treatment of the bandlimit of the functions (based on Karras et al. (2021)), the convolutional neural operator (Raonić et al., 2023) also uses up- and downsampling of the input and output functions to apply a U-Net architecture.
@@ 3.1 Motivation: Convolutional Layer
$z_{i}=h\left(i-1-\tfrac{S-1}{2}\right),$ (2)
@@ 3.2 Differential Layer
In this section, we construct a layer that converges to a differential operator when the width $h$ of the discretization $D^{h}$ tends to zero. To prevent the operator from collapsing to a pointwise operation, we constrain and rescale the values of the kernel (according to the discretization width). Note that without rescaling, we would again recover a pointwise operator; however, due to the rescaling by a factor of $\frac{1}{h}$, the limit might not exist without constraining the values of the kernel. The next proposition shows that it is sufficient to subtract the average kernel value. See Section A.1 for a more general statement and a corresponding proof.
@@
Proposition 3.1 shows that we can learn different directional derivatives using an appropriate adaptation of standard convolutional layers as in Section 3.1. Specifically, we center the kernel $K$ by subtracting its mean $\bar{K}$ and scale the result by the reciprocal resolution $\frac{1}{h}$. In Figure 6, we empirically evaluate the convergence in resolution to a unique differnetial operator on a simple example.
@@ 3.3 Integral Kernel Layers
Instead of rescaling the kernel, we can also adapt the size of the kernel such that the receptive field stays the same, i.e., is independent of the resolution $h$. In this section, we begin at the general graph neural operator and demonstrate how (considering the desirable properties of translation equivariance and efficiency) we arrive at local convolutional layers with adaptive kernel sizes using a fixed number of parameters.
Graph neural operator (GNO):
@@
Fourier Neural Operator (FNO):
@@
The remaining question is centered around the parametrization of the kernel. One could draw inspiration from works on hypernetworks in computer vision and parameterize it using a neural network (Shocher et al., 2020). However, this is significantly more costly than a standard convolution. Another idea is to interpolate a fixed-sized kernel, e.g., using sinc or bi-linear interpolation. A more general version of the latter, based on a learnable linear combination of hat functions, is also known as discrete-continuous (DISCO) convolution and has shown to be effective in different applications (Ocampo et al., 2022). We note that this formulation is closest to the original convolutional layer typically used in computer vision. However, in the next section, we see that it also allows us to use unstructured meshes and formulate the operation on more general domains $D\subset\mathbb{R}^{d}$.
@@ 3.4 General discrete-continuous convolutions
where $g,x\in G$ and $\mathrm{d}\mu(x)$ is the invariant Haar measure on $G$. This formulation presents us with the challenge of being non-trivial to discretize. The framework of DISCO convolutions (Ocampo et al., 2022) achieves this by approximating the integral with a quadrature rule while evaluating the group action continuously:
@@ 3.5 Local neural operator architecture
To design our neural operator, we want to combine pointwise, local, and global operations. To this end, we take an FNO (or SFNO for spherical problems) as a starting point, which already features global operations in Fourier space and pointwise operations using its residual connections. Then, we augment the operators by incorporating our proposed local convolutions from the previous section as additional branches in the respective layers (see Figure 2). These four branches are summed pointwise within each local neural operator layer. The resulting architecture can be trained end-to-end with any standard operator learning training procedure and loss.